mathslover (mathslover):

Logarithm Question.... If y = \(a^{\left(\cfrac{1}{1-\log_a x}\right)} \) and z = \(a^{\left(\cfrac{1}{1-\log_a y}\right)}\) then x = ?

mathslover (mathslover):

If y = \(a^{\left(\cfrac{1}{1-\log_a x}\right)} \) and z = \(a^{\left(\cfrac{1}{1-\log_a y}\right)}\)

mathslover (mathslover):

@Kainui and @iambatman

mathslover (mathslover):

This is what I have done yet : In order to get rid of the powers, I took log both sides : \(\log y = \cfrac{1}{1-\log_a x} \log a\) and \(\log z = \cfrac{1}{1-\log_a y} \log a\)

mathslover (mathslover):

Okay, let us try to simplify it more. \(\log y = \left(\cfrac{1}{1-\log_a x} \right)\left(\cfrac{1 }{\log x}\right)\) \(\log x \times \left(1-\log_a x\right) = \cfrac{1}{\log y}\) \(\log x \times \left(1-\cfrac{\log a}{\log x }\right) = \cfrac{1}{\log y}\)

mathslover (mathslover):

Seems like, I have made it more complicated.

mathslover (mathslover):

This is what I have got after a lot of work : \(\log_a \left(\cfrac{a}{x}\right) = \log_a \left(\cfrac{a}{y}\right) \times \left(\cfrac{\log y}{\log z}\right)\)

mathslover (mathslover):

and making the bases same for everything : \(\log_a \left(\cfrac{a}{x}\right) = \log_a \left(\cfrac{a}{y}\right) \times \left(\cfrac{\log_a z}{\log_a y}\right) \)

mathslover (mathslover):

\(\cfrac{a}{x} = \left(\cfrac{a}{y}\right)^{\left(\cfrac{\log_a z}{\log_a y}\right)}\)

OpenStudy (zzr0ck3r):

\(\log_a(z)=\frac{1}{1-\log_a(y)}\\1-\frac{1}{\log_a(z)}=\log_a(y)\\y=a^{\frac{\log_a(z-1)}{\log_a(z)}}\) now \(y=a^{\frac{1}{1-\log_a(x)}}\implies a^{\frac{\log_a(z-1)}{\log_a(z)}}=a^{\frac{1}{1-\log_a(x)}}\implies \frac{\log_a(z-1)}{\log_a(z)}=\frac{1}{1-\log_a(x)}\implies \\x=a^{\frac{\log_a(z)-\log_a(z-1)}{\log_a{z-1}}}\)

OpenStudy (zzr0ck3r):

I think...

mathslover (mathslover):

What you have written is though, not visible ot me, but yes, I'm getting what you did.

OpenStudy (zzr0ck3r):

brb

OpenStudy (kainui):

I'm getting after a lot of disgusting looking algebra that if we use the same form of function, x(z) has the same form as y(x) and z(y) shown above. \[\Large x=a^{\frac{1}{1-\log_az}}\]

mathslover (mathslover):

Well, may be these options help us : \(\boxed{\\ \text{(A)} \quad a^{\cfrac{1}{1+\log_a z}} \\ \text{(B)} \quad a^{\cfrac{1}{2+ \log_a z}} \\ \text{(C)} \quad a^{\cfrac{1}{1-\log_a z}} \\ \text{(D)} \quad \text{None of these} \\ }\)

mathslover (mathslover):

@Kainui - Yeah your answer seems to be right, as it is in the options.

OpenStudy (kainui):

I'm getting c, and here's my algebra for checking give me a moment...

OpenStudy (kainui):

I solved in general, \[\Large h=a^{\frac{1}{1-\log k}} \]\[\Large \log h=\frac{1}{1-\log k}\]\[\Large 1-\log k=\frac{1}{\log h}\]\[\Large 1-\frac{1}{\log h}= \log k\]\[\Large k=a^{ 1-\frac{1}{\log h}}\] Now since h(k) is the same form as both y(x) and z(y) we can plug in each equation to itself... One sec...

OpenStudy (kainui):

We now have found k(h) so we really have x(y) and y(z) now so we simply find x(y(z)) since that's what all our answer choices are.

OpenStudy (kainui):

\[\Huge x(y(z))=a^{ 1-\frac{1}{\log (a^{ 1-\frac{1}{\log h}})}}\]The log and a part in the exponent cancel out. \[\Huge x=a^{ 1-\frac{1}{ 1-\frac{1}{\log h}}}\] From here we simplify the fractions \[\Huge x=a^{ 1-\frac{1}{ \frac{\log h -1}{\log h}}}\]\[\Huge x=a^{ 1-\frac{\log h}{ \log h -1}}\]\[\Large x=a^{ \frac{\log h -1 -\log h}{ \log h -1}}\]\[\Large x=a^{ \frac{-1}{ \log h -1}}=a^{ \frac{1}{ 1-\log h }}\] But I can't help but feel like there's a better way of doing this one. Also I just realized all those h's should be z's. Whatever.

mathslover (mathslover):

Well, yeah, I think this is one of the ways of solving this quesiton,,,, but there must be an easier way to get to the answer.

OpenStudy (kainui):

Obviously there's a symmetry going on here, but it beats me how it works. Really cool problem.

mathslover (mathslover):

This is what we can do : \((1-\log_a x)(\log y) = (1-\log_a y)(\log z)\)

mathslover (mathslover):

I think I got it.

OpenStudy (kainui):

What is it? I think there's a really nice solution to this that's really fast.

OpenStudy (kainui):

This might be better: h(k) has the form of y(x) and z(y) k(h) has the form x(y) and y(z). So we can set y(x)=y(z) if we just play around with the general equation. \[\Large h=a^{1-\frac{1}{\log k}}\] Ok so skipping the little algebra I get to these two equations from this one. \[\Large \log h=1-\frac{1}{\log k}\]\[\Large \log k=\frac{1}{1-\log h}\] Above when I said y(x)=y(z) this also means log[y(x)]=log[y(z)] So let's do that. \[1-\frac{1}{\log x}=\frac{1}{1-\log z}\] Solve for x. I don't know if this is really that much better hmm... I feel like there's a really satisfying way to recognize that y(x), z(y) and x(z) are all identical forms without having to do anything at all.

mathslover (mathslover):

Sorry, my net stopped working.

mathslover (mathslover):

Yeah, I agree Kainui that there should be much satisfying way to do this.